x + 3y = 7
-x + 4y = 7

Answer by checkley77(12844)

x + 3y = 7
-x + 4y = 7 add
-----------------
7y=14
y=14/7
y=2 ans.
x+3*2=7
x+6=7
x=7-6
x=1 ans.
Proof:
-1+4*2=7
-1+8=7
7=7